# Within the Dynamic Context: Inertia-aware 3D Human Modeling with Pose Sequence

### v1

@@ 1.1 Localized Dynamic Context Encoding:
Kinematically-guided Spatial Dependency: For a given context sequence $S_{\Delta{\bm{p}}}$ and an input coordinate $\bm{x}$, we begin by processing each $\Delta{\bm{p}}\in\mathbb{R}^{3\times K+3}$ independently within the sequence. Here we leverage a reasonable physical assumption: the impact of a rotating joint is mainly restricted to the kinematic chains to which it belongs. For instance, the movement of the hips generally has minimal impact on the appearance of the shoulders. Therefore, the motion and appearance of a point $\bm{x}$ is expected to mainly depend on the joints within the same kinematic chains. With an input coordinate $\bm{x}$ and its associated blending weights $W_{c}(\bm{x})\in\mathbb{R}^{K}$, we first identify its nearest joint $k=\operatorname*{arg\,max}W_{c}(\bm{x})$ to determine the kinematic chains $\bm{x}$ lies on. We use the kinematic chains predefined in the skeleton model used by SMPL. Then we gather all joints from the related kinematic chains. Last, we apply a binary mask to each $\Delta{\bm{p}}$ to filter out the unrelated joints that fall outside the kinematic chains. We denote masked conditions as $\hat{\Delta}{\bm{p}}$.
Spatial Encoding: We then flatten the masked spatial condition $\Delta{\bm{p}}$ to an 1D vector and project it to a $16$-dim vector by an one-layered MLP with ReLU activation. We pass the vector of each time step through the same MLP, resulting in a sequence of $16$-dimensional vectors with a length of $L$.
@@ 1.2 Other Modules:
Canonical Volume: We implement the canonical volume $F_{c}$ as a tri-plane encoder [6], followed by an MLP with a color branch and a sigma branch. As for the tri-plane grids, we use multi-scale planes with $4$ different resolutions at $64^{2}$, $128^{2}$, $256^{2}$, and $512^{2}$. The per-plane and per-scale features are multiplied across planes and concatenated across scales. We set each per-plane and per-scale dimension as $32$ so the final encoded feature has dimension $128$. The Color MLP has 2 hidden layers with a channel dimension as $256$ and the sigma branch has a single hidden layer with a channel dimension of $256$.
@@
Pose-sequence Related Input: For the delta-pose sequence condition $S_{\Delta{\bm{p}}}=\{\Delta{\bm{p}_{i-(L-1)s}},\ldots,\Delta{\bm{p}_{i}}\}$. We set $L=6$, $s=25$. So the context covers the motion history of 150 frames forward. We embed the angles rotating the joints from the last step to the current step and represent them in the axis-angle form. As the pose difference between two consecutive frames can be small and noisy, we take a stride $s_{d}=25$ backwards and compute the delta as $\Delta{\bm{p}_{i}}=\delta(\bm{p}_{i},\bm{p}_{i-s_{d}})$. Additionally, we include a 3-dimensional global translation vector and concatenate it with the joint rotation vectors, yielding an $L$-length sequence of delta vectors with a dimension of $33K+$ 3.
@@ 3 I3D-Human Dataset
Figure 9: Visualization of our I3D-Human dataset. Our dataset focuses on capturing variations in clothing appearance under approximately identical poses. We will release the dataset for research purposes.
@@ 4 Results on ZJU-MoCap
We show the qualitative comparison on the ZJU-MoCap dataset in Fig. 13. In novel-view test set, our method can capture more details on the clothing wrinkles. Although NeuralBody [28] attains the highest PSNR and SSIM scores in novel views, it suffers from significant blurriness. The novel pose renderings generated by our method are quite comparable to those produced by HumanNeRF [40] and 3DGS-Avatar [33].
@@
PSNR↑ SSIM↑ LPIPS↓ DME↓ PSNR↑ SSIM↑ LPIPS↓ DME↓
@@
(a) Inputting $S_{\Delta{\bm{p}}}$ outperforms $S_{\bm{p}}$.
@@
PSNR↑ SSIM↑ LPIPS↓ DME↓ PSNR↑ SSIM↑ LPIPS↓ DME↓
@@
PSNR↑ SSIM↑ LPIPS↓ DME↓ PSNR↑ SSIM↑ LPIPS↓ DME↓
@@ 5.1 Pose Sequence Condition
We study the choice of the input dynamic context sequence, including the following factors.
Pose Sequence or Delta Pose Sequence: As for each element in the input context sequence, instead of the pose parameter at each step, we use the pose difference between two consecutive steps including the angular and translational velocities. We call the resulting sequence as the delta pose sequence $S_{\Delta{\bm{p}}}$. We compare inputting pose sequence $S_{\bm{p}}$ and delta pose sequence $S_{\Delta{\bm{p}}}$ and show the results on I3D-Human in Tab. 6(a). Inputting $S_{\Delta{\bm{p}}}$ achieves better performance on both novel view and novel pose, thanks to the reduction of the input complexity and comprehensive inclusion of all dynamic contexts.
@@
Length and Step: We further ablate on different choices of delta pose sequence length $L_{d}$, sequence step $s$, and delta step $s_{d}$. Due to similar physical meanings of $s$ and $s_{d}$, we keep them the same in our experiments. The ablation results can be found in Tab. 6(c). Short sequence ($L_{d}=2$) lacks sufficient dynamic contexts, leading to a decrease in performance on both novel view and novel pose. On the other hand, long sequence ($L_{d}=12$) contains overly complex input, which leads to comparable performance on novel view but worse overfitting performance on novel pose. Reducing $s$ and $s_{d}$ ($s\,\&\,s_{d}=12$) similarly leads to a loss of dynamic contexts, resulting in deteriorated performance on both novel view and novel pose. We rigorously select the most suitable parameter combinations to encompass an adequate amount of dynamic contexts while mitigating overfitting induced by complex inputs.
@@ 5.2 The effectiveness of Localized Spatial Dependency
Figure 10: Localized spatial dependency enhances generalization for novel poses composed of observed sub-poses. For instance, consider a training pose (left) where we generate a novel pose by substituting the pose parameters of the upper limbs with those from another training frame (right). This novel combination of pose parameters challenges models with global dependency on pose conditions, leading to inaccuracies in rendering dynamic-related details on the pant legs. In contrast, our method with local pose dependency can still capture the appearance variation.
To showcase the advantage of localized spatial dependency, we examine novel pose combinations by recombining sub-poses observed in the training set. For instance, as illustrated in Fig. 10, suppose we have a video sequence where the character waves her arms and legs in opposite directions. We create a novel pose by substituting the sub-pose of the upper limbs with the one from another time-step. Thanks to the observed sub-poses and the limited interdependence between the upper and lower limbs, it is expected that they should be rendered akin to their appearance in the training frames under this novel pose combination. We conduct a comparison between renderings produced by the model with local spatial dependency (our default option) and the other one with global spatial dependency. The results demonstrate that our method, with local dependency, effectively captures the motion of the pant leg when the character swings her legs up or down. In contrast, the model with global dependency fails to accurately render the legs under the novel pose combination given the encoded global context condition is unseen.
@@ 6 Limitations
Firstly, similar to previous implicit avatars, Dyco encounters challenges when animating out-of-distribution novel poses. Although we propose localized spatial dependency to help generalization towards unseen poses composed of seen sub-poses, it still struggles in predicting appearance and geometry variation under completely novel poses and inpainting unseen body parts. We provide an illustrative example from the ZJU-MoCap novel pose test set in Fig. 11, where the character lifts his elbow, revealing the underarm region that was not observed during training. It’s important to note that other methods also fail to render the underarm plausibly.

### v2

@@ 6 Conclusion
Acknowledgements
This work is supported by the Shanghai Artificial intelligence Laboratory, in part by JSPS KAKENHI Grant Numbers 24K22318, 22H00529, 20H05951, JST-Mirai Program JPMJMI23G1.
@@ 1.1 Localized Dynamic Context Encoding:
Kinematically-guided Spatial Dependency: For a given context sequence $S_{\Delta{\bm{p}}}$ and an input coordinate $\bm{x}$, we begin by processing each $\Delta{\bm{p}}\in\mathbb{R}^{3\times K+3}$ independently within the sequence. Here we leverage a reasonable physical assumption: the impact of a rotating joint is mainly restricted to the kinematic chains to which it belongs. For instance, the movement of the hips generally has minimal impact on the appearance of the shoulders. Therefore, the motion and appearance of a point $\bm{x}$ is expected to mainly depend on the joints within the same kinematic chains. With an input coordinate $\bm{x}$ and its associated blending weights $W_{c}(\bm{x})\in\mathbb{R}^{K}$, we first identify its nearest joint $k=\operatorname*{arg\,max}W_{c}(\bm{x})$ to determine the kinematic chains $\bm{x}$ lies on. We use the kinematic chains predefined in the skeleton model used by SMPL. Then we gather all joints from the related kinematic chains. Last, we apply a binary mask to each $\Delta{\bm{p}}$ to filter out the unrelated joints that fall outside the kinematic chains. We denote masked conditions as $\hat{\Delta}{\bm{p}}$. We describe the effectiveness of this design in Sec. 5.2.
Spatial Encoding: We then flatten the masked spatial condition $\hat{\Delta}{\bm{p}}$ to an 1D vector and project it to a $16$-dim vector by an one-layered MLP with ReLU activation. We pass the vector of each time step through the same MLP, resulting in a sequence of $16$-dimensional vectors with a length of $L$.
@@ 1.2 Other Modules:
Canonical Volume: We implement the canonical volume $F_{c}$ as a tri-plane encoder [6], followed by an MLP with a color branch and a density branch. As for the tri-plane grids, we use multi-scale planes with $4$ different resolutions at $64^{2}$, $128^{2}$, $256^{2}$, and $512^{2}$. The per-plane and per-scale features are multiplied across planes and concatenated across scales. We set each per-plane and per-scale dimension as $32$ so the final encoded feature has dimension $128$. The Color MLP has 2 hidden layers with a channel dimension as $256$ and the density branch has a single hidden layer with a channel dimension of $256$.
@@
Pose-sequence Related Input: For the delta-pose sequence condition $S_{\Delta{\bm{p}}}=\{\Delta{\bm{p}_{i-(L-1)s}},\ldots,\Delta{\bm{p}_{i}}\}$. We set $L=6$, $s=25$. So the context covers the motion history of 150 frames forward. We embed the angles rotating the joints from the last step to the current step and represent them in the axis-angle form. As the pose difference between two consecutive frames can be small and noisy, we take a stride $s_{d}=25$ backwards and compute the delta as $\Delta{\bm{p}_{i}}=\delta(\bm{p}_{i},\bm{p}_{i-s_{d}})$. Additionally, we include a 3-dimensional global translation vector and concatenate it with the joint rotation vectors, yielding an $L$-length sequence of delta vectors with a dimension of $33K+$ 3, where $K$ is the number of joints.
@@ 3 I3D-Human Dataset
Figure 9: Visualization of our I3D-Human dataset. Our dataset focuses on capturing variations in clothing appearance under approximately identical poses. The dataset is released for research purposes.
@@ 4 Results on ZJU-MoCap
Besides the quantitative comparison in Tab. 2 of the main paper, we show the qualitative comparison on the ZJU-MoCap dataset in Fig. 13. In novel-view test set, our method can capture more details on the clothing wrinkles. Although NeuralBody [28] attains the highest PSNR and SSIM scores in novel views, it suffers from significant blurriness. The novel pose renderings generated by our method are quite comparable to those produced by HumanNeRF [40] and 3DGS-Avatar [33].
@@
[…]
@@
(a) Inputting $S_{\Delta{\bm{p}}}$ outperforms $S_{\bm{p}}$. $\textrm{LPIPS}^{*}=\textrm{LPIPS}\times 10^{3}$
@@
[…]
@@
[…]
@@ 5.1 Pose Sequence Condition
We ablate the design of the dynamic context sequence input, including the following factors.
Pose Sequence or Delta Pose Sequence: As for each element in the input context sequence, instead of the pose parameter at each step, we use the pose difference between two consecutive steps including the angular and translational velocities. We call the resulting sequence as the delta pose sequence $S_{\Delta{\bm{p}}}$. We compare inputting pose sequence $S_{\bm{p}}$ and delta pose sequence $S_{\Delta{\bm{p}}}$, and show the results on I3D-Human in Tab. 6(a). Inputting $S_{\Delta{\bm{p}}}$ achieves better performance on both novel view and novel pose, thanks to the reduction of the input complexity and comprehensive inclusion of all dynamic contexts.
@@
Length and Step: We further ablate on different choices of delta pose sequence length $L_{d}$, sequence step $s$, and delta step $s_{d}$. We set $s$ and $s_{d}$ the same in our experiments. The ablation results can be found in Tab. 6(c). Short sequence ($L_{d}=2$) lacks sufficient dynamic contexts, leading to a decrease in performance on both novel view and novel pose. On the other hand, long sequence ($L_{d}=12$) contains overly complex input, which leads to comparable performance on novel view but worse overfitting performance on novel pose. Reducing $s$ and $s_{d}$ ($s\,\&\,s_{d}=12$) similarly leads to a loss of dynamic contexts, resulting in deteriorated performance on both novel view and novel pose. We rigorously select the most suitable parameter combinations to encompass an adequate amount of dynamic contexts while mitigating overfitting induced by complex inputs.
@@ 5.2 The effectiveness of Localized Spatial Dependency
Figure 10: Localized spatial dependency enhances generalization for novel poses composed of observed sub-poses. Consider a training pose (left) from which we create a novel pose (right) by substituting the upper limbs’ pose with that from another training frame. This novel combination of sub-poses challenges models with global dependency, leading to inaccuracies in rendering dynamic-related details on the pant legs. In contrast, our method with local pose dependency can still capture the appearance variation.
We quantitatively measure the impact of local or global spatial dependency in Tab. 3 of the main text, where the last two rows indicate that localization improves three conventional metrics on the test set. However, we observe that localization does not enhance DME. This may be due to: (1) DME computation relying on optical flow estimation, which may contain errors, and (2) our dataset’s lack of diverse pose combinations, which limits the ability to highlight the importance of avoiding spurious inter-body dependency. To address this limitation and prove the effectiveness of localized spatial dependency, we examine novel pose combinations by recombining sub-poses observed in the training set. As illustrated in Fig. 10, suppose we have a video sequence where the character waves her arms and legs in opposite directions. We create a novel pose by substituting the sub-pose of the upper limbs with the one from another time-step. Thanks to the observed sub-poses and the limited interdependence between the upper and lower limbs, it is expected that they should be rendered akin to their appearance in the training frames under this novel pose combination. We compare renderings by two models with local and global spatial dependency. We show that the one with local dependency effectively captures the local motion of the pant leg, while the other one with global dependency fails to accurately render the legs under the novel pose combination as the encoded global context condition is unseen.
@@ 6 Limitations
Firstly, similar to previous implicit avatars, Dyco encounters challenges when animating out-of-distribution novel poses. Although we propose localized spatial dependency to help generalization towards unseen poses composed of seen sub-poses, as shown in Fig. 10, it still struggles in predicting appearance and geometry variation under completely novel poses and inpainting unseen body parts. We provide an illustrative example from the ZJU-MoCap novel pose test set in Fig. 11, where the character lifts his elbow, revealing the underarm region that was not observed during training. It’s important to note that other methods also fail to render the underarm plausibly.
@@
Lastly, the proposed dynamic motion error (DME) metric relies on the accurate estimation of 2D optical flow. However, we find that off-the-shelf optical flow estimators, such as RAFT [37], are not perfect. While DME can generally measure how accurately the rendering exhibits motion dynamics, as shown in Fig. 8, it may not strictly correlate with performance at a finer scale, as shown in some of our ablation results. Research progress in 3D point tracking might shed light on the measurement of motion dynamics.